OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 + 12x + 30 8x – y = 10

OpenStudy (anonymous):

I won't give you the answer bgut I'll help you solve it. First, isolate your Y's.

hero (hero):

HInt: 8x - 10 = x^2 + 12x + 30

OpenStudy (anonymous):

Well, I guess someone will do it for you.

OpenStudy (anonymous):

\[0=x^2+4x+20\]

hero (hero):

It was a hint. She still has to find the solutions

OpenStudy (anonymous):

WolframAlpha to the rescue!!

hero (hero):

Wow, really

OpenStudy (anonymous):

and its giving weird answers .-.

OpenStudy (anonymous):

I mean x² + 4x + 40 = 0

OpenStudy (anonymous):

yea thats what i put

OpenStudy (anonymous):

its giving complex solutions

hero (hero):

@Chlorophyll, I thought you were trying to give someone plus 40 likes

hero (hero):

yummydum, do b^2-4ac . If it is negative, solutions will be complex

OpenStudy (anonymous):

@Hero Not anymore ;)

OpenStudy (anonymous):

its negative so does that mean there are no real solutions

hero (hero):

Yup

OpenStudy (anonymous):

So do you need to find complex roots?

OpenStudy (anonymous):

no its multiple choice n i picked no real solutions :)

OpenStudy (anonymous):

Look like happy ending :D

OpenStudy (anonymous):

What is the solution to the following system of equations? y = x2 + 10x + 11 y = x2 + x – 7 (–2, –5) (2, –5) (–2, 5) (2, 5

OpenStudy (anonymous):

lol @Chlorophyll if only

OpenStudy (anonymous):

@Hero i know its A or C

OpenStudy (anonymous):

Open the new post for each of question! Anyway, equal the 2 equations!

OpenStudy (anonymous):

its A?

OpenStudy (anonymous):

it is ^_^

OpenStudy (anonymous):

Correct, A!

hero (hero):

@yummydum, if you do the math, you'll know there's only one possible solution.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

i did-ed it all by myself ^_~